The two taylor girls were born in consecutive, even years. If the product of their age is 224, how old is each girl?
Answer by checkley77(12844)

x(x+2)=224
x^2+2x-224=0
(x-14)(x+16)=0
x-14=0
x=14 for the youngest girl.
14+2=16 for the older girl.
Proof:
14*16=224
224=224